# Fanxu Meng1,2,3, Zhaohui Wang1, Muhan Zhang1,3

### v3

@@ Abstract
To parameter-efficiently fine-tune (PEFT) large language models (LLMs), the low-rank adaptation (LoRA) method approximates the model changes $\Delta W\in\mathbb{R}^{m\times n}$ through the product of two matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, $A$ is initialized with Gaussian noise, and $B$ with zeros. LoRA freezes the original model $W$ and updates the “Noise & Zero” adapter, which may lead to slow convergence. To overcome this limitation, we introduce Principal Singular values and Singular vectors Adaptation (PiSSA). PiSSA shares the same architecture as LoRA, but initializes the adaptor matrices $A$ and $B$ with the principal components of the original matrix $W$, and put the remaining components into a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ which is frozen during fine-tuning. Compared to LoRA, PiSSA updates the principal components while freezing the “residual” parts, allowing faster convergence and enhanced performance. Comparative experiments of PiSSA and LoRA across 12 different models, ranging from 184M to 70B, encompassing 5 NLG and 8 NLU tasks, reveal that PiSSA consistently outperforms LoRA under identical experimental setups. On the GSM8K benchmark, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, surpassing LoRA’s 67.7% by 5.16%. Due to the same architecture, PiSSA is also compatible with quantization to further reduce the memory requirement of fine-tuning. Compared to QLoRA, QPiSSA (PiSSA with 4-bit quantization) exhibits smaller quantization errors in the initial stages. Fine-tuning LLaMA-3-70B on GSM8K, QPiSSA attains an accuracy of 86.05%, exceeding the performances of QLoRA at 81.73%. Leveraging a fast SVD technique, PiSSA can be initialized in only a few seconds, presenting a negligible cost for transitioning from LoRA to PiSSA.
@@
Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA. In this visualization, blue modules represent parts of the model where parameters are frozen during training, while orange modules indicate components that require updates. QLoRA quantizes the pretrained matrix in LoRA to 4-bit, whereas QPiSSA quantizes the residual matrix in PiSSA.
@@
Figure 2: We illustrate the two key advantages of PiSSA: converging faster and better, and reducing quantization error. In the left figure, we use a toy example to show PiSSA’s faster convergence, where we first train a two-layer MLP classifying odd numbers of MNIST, and then fine-tune the model on even numbers. PiSSA finds the right direction more quickly and achieves a lower loss with the same number of steps. In the right figure, we show the quantization error reduction of PiSSA compared to directly quantizing the base model. PiSSA can reduce more quantization error than the baselines.
1 Introduction
Fine-tuning large language models (LLMs) is a highly effective technique for boosting their capabilities in various tasks [1, 2, 3, 4], ensuring models to follow instructions [5, 6, 7], and instilling models with desirable behaviors while eliminating undesirable ones [8, 9]. However, the fine-tuning process for very large models is accompanied by prohibitive costs. For example, regular 16-bit fine-tuning of a LLaMA 65B parameter model requires over 780 GB of GPU memory [10], and the VRAM consumption for training GPT-3 175B reaches 1.2TB [11]. Consequently, various parameter-efficient fine-tuning (PEFT) [12, 13] methods have been proposed to reduce the number of parameters and memory usage required for fine-tuning. Due to the ability to maintain the performance of full fine-tuning without adding additional inference latency, Low-Rank Adaptation (LoRA) [11] has emerged as a popular PEFT method.
LoRA [11] hypothesizes that the modifications to parameter matrices during fine-tuning exhibit low-rank properties. As depicted in Figure 1(b), for a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA substitutes the updates with a low-rank decomposition $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, and the rank $r\ll\text{min}(m,n)$. For $Y=XW$, the modified forward pass is as follows:
$Y=X(W+\Delta W)=X(W+AB),$ (1)
where $X\in\mathbb{R}^{b\times m}$, $Y\in\mathbb{R}^{b\times n}$, and $b$ represents the batch size of input data. A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially. LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix $W$, instead optimizing the injected, significantly smaller low-rank matrices $A,B$. Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times [11]. Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks. By integrating the quantization of pre-trained matrices $W$, LoRA also enables reducing the average memory requirements by 16 times [10]. Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not significantly degrade the performance of LoRA.
According to Equation 1, the gradients of A and B are $\frac{\partial L}{\partial A}=X^{T}\left(\frac{\partial L}{\partial Y}\right)B^{T}$ and $\frac{\partial L}{\partial B}=A^{T}X^{T}\left(\frac{\partial L}{\partial Y}\right)$. Compared to full fine-tuning, using LoRA initially does not change the output $Y$ for the same input $X$, so the gradient magnitude is primarily determined by the values of $A$ and $B$. Since $A$ and $B$ are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process. We also observe this phenomenon empirically, where LoRA often wastes much time around the initial point.
@@ 1 Introduction
Since the principal components $W^{pri}$ are preserved in the adapter and are of full precision, another benefit of PiSSA is that when applying quantization to the frozen part $W^{res}$, we can significantly reduce the quantization error compared to QLoRA (which quantizes the whole $W$), as illustrated in Figure 2(b). Therefore, PiSSA is perfectly compatible with quantization, making it a plug-and-play substitution for LoRA.
@@ 2 Related Works
The vast complexity and computational needs of large language models (LLMs) with billions of parameters present significant hurdles in adapting them for specific downstream tasks. Parameter Efficient Fine-Tuning (PEFT) [12, 13] emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning. PEFT encompasses strategies like partial fine-tuning [14, 15, 16, 17, 18, 19, 20, 21], soft prompt fine-tuning  [22, 23, 24, 25, 26, 27, 28], non-linear adapter fine-tuning [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and low rank adapter based fine-tuning [39, 40, 11, 41].
LoRA [11] inject trainable adapters to the linear layers. After fine-tuning, these adaptations can be re-parameterized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning. Following LoRA, AdaLoRA [42, 41, 43] dynamically learns the rank size needed for LoRA in each layer of the model. DeltaLoRA [44, 45] updates the original weights of the model using parameters from adapter layers, enhancing LoRA’s representational capacity. LoSparse [46] incorporates LoRA to prevent pruning from eliminating too many expressive neurons. DoRA [47] introduces a magnitude component to learn the scale of $\Delta W$ while utilizing the original AB as a direction component of $\Delta W$. Unlike LoRA and its successors, which focus on learning low-rank approximations of weight updates, our PiSSA approach directly tunes the essential but low-rank parts of the model while keeping the noisier, high-rank, and nonessential parts frozen. Although our approach differs in philosophy from LoRA, it shares most of LoRA’s structural benefits and can be extended by these methods to enhance its performance.
QLoRA [10] integrates LoRA with 4-bit NormalFloat (NF4) quantization, along with Double Quantization and Paged Optimizers, enabling the fine-tuning of a 65B parameter model on a single 48GB GPU while preserving the performance of full 16-bit fine-tuning tasks. QA-LoRA [48] introduces group-wise operators to increase the degree of freedom in low-bit quantization. LoftQ [49] reduces quantization error by decomposing the quantization error matrix of QLoRA and retaining the principal components with an adapter. Our PiSSA approach can also be combined with quantization techniques, and we have found that PiSSA significantly reduces quantization error compared to QLoRA and LoftQ.
@@ 3 PiSSA: Principal Singular Values and Singular Vectors Adaptation
Similar to LoRA, the gradients of $A$ and $B$ are also given by $\frac{\partial L}{\partial A}=X^{T}\left(\frac{\partial L}{\partial Y}\right)B^{T}$ and $\frac{\partial L}{\partial B}=A^{T}X^{T}\left(\frac{\partial L}{\partial Y}\right)$. Since elements of $\mathbf{s}_{[:r]}$ $\gg$ elements of $\mathbf{s}_{[r:]}$, the trainable adapter $W^{pri}=AB$ contains the most essential directions of $W$. In the ideal case, training $AB$ mirrors the process of fine-tuning the entire model despite using fewer parameters. The ability to directly fine-tune the most essential part of a model enables PiSSA to converge faster and better. In contrast, LoRA initializes the adapters $A$ and $B$ with Gaussian noise and zeros while keeping $W$ frozen. Consequently, the gradients are small or in random directions during the early stages of fine-tuning, possibly introducing much waste of gradient descent steps. Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits. These include but are not limited to the capability of fine-tuning a model with a reduced number of trainable parameters, quantizing the residual model to decrease memory consumption during forward propagation in training, and easy deployment. The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model. Employing the Fast SVD technique [50] allowed PiSSA to finish initialization in several seconds (Appendix B), which is a negligible cost.
@@
4 QPiSSA: PiSSA with Quantization
Quantization divides the value range of a matrix into several continuous regions, and maps all values falling inside a region into the same “quantized” value. It is an effective technique to reduce the memory consumption of forward propagation, but breaks down during backpropagation. At the same time, LoRA greatly reduces the backward memory requirement, making it highly suitable to use LoRA and quantization together, where the base model is quantized for memory-efficient forward propagation, and the LoRA adaptors are kept in full precision for accurate backward parameter updates. One representative previous work, QLoRA, quantizes the base model to Normal Float 4-bit (NF4) and initializes the full-precision $A$ and $B$ with Gaussian-Zero initialization. Therefore, the overall error is given by:
@@ 4 QPiSSA: PiSSA with Quantization
where $||M||_{*}$ denotes the nuclear norm (also known as the trace norm)  [51], defined as:
@@
Since the residual model has removed the large-singular-value components, $W^{res}$ has a narrower distribution than that of $W$, as can be seen in Figures 3(a) and 3(b) (comparing the singular value distributions of $W$ and $W^{res}$), as well as Figures 3(c) and 3(f) (comparing the value distributions of $W$ and $W^{res}$), which is beneficial for reducing the quantization error. Furthermore, since NF4 is optimized for normally distributed data [10], we fit a Gaussian for the values of $W$ and $W^{res}$, respectively. As can be seen in Figures 3(c) and 3(f), $W^{res}$ is much more Gaussian-like and has a smaller standard deviation, making it more suitable to apply NF4 to $W^{res}$ instead of $W$. Both the above lead QPiSSA to achieve a significantly lower quantization error than QLoRA, shown in Figures 3(d) and 3(e).
@@
Figure 3: Figures (a), (b), (d) and (e) visualize the singular values in descending order for $W$, $W^{res}$, $W-nf4(W)$ and $W^{res}-nf4(W^{res})$ of LLaMA 2-7B’s “layers[0].self_attn.q_proj” matrix. Figures (c) and (f) show the data value distributions of $W$ and $W^{res}$.

### v4

# Fanxu Meng1,2, Zhaohui Wang1, Muhan Zhang1,2

@@ Abstract
To parameter-efficiently fine-tune (PEFT) large language models (LLMs), the low-rank adaptation (LoRA) method approximates the model changes $\Delta W\in\mathbb{R}^{m\times n}$ through the product of two matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, $A$ is initialized with Gaussian noise, and $B$ with zeros. LoRA freezes the original model $W$ and updates the “Noise & Zero” adapter, which may lead to slow convergence. To overcome this limitation, we introduce Principal Singular values and Singular vectors Adaptation (PiSSA). PiSSA shares the same architecture as LoRA, but initializes the adaptor matrices $A$ and $B$ with the principal components of the original matrix $W$, and put the remaining components into a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ which is frozen during fine-tuning. Compared to LoRA, PiSSA updates the principal components while freezing the “residual” parts, allowing faster convergence and enhanced performance. Comparative experiments of PiSSA and LoRA across 11 different models, ranging from 184M to 70B, encompassing 5 NLG and 8 NLU tasks, reveal that PiSSA consistently outperforms LoRA under identical experimental setups. On the GSM8K benchmark, Gemma-7B fine-tuned with PiSSA achieves an accuracy of 77.7%, surpassing LoRA’s 74.53% by 3.25%. Due to the same architecture, PiSSA is also compatible with quantization to further reduce the memory requirement of fine-tuning. Compared to QLoRA, QPiSSA (PiSSA with 4-bit quantization) exhibits smaller quantization errors in the initial stages. Fine-tuning LLaMA-3-70B on GSM8K, QPiSSA attains an accuracy of 86.05%, exceeding the performance of QLoRA at 81.73%. Leveraging a fast SVD technique, PiSSA can be initialized in only a few seconds, presenting a negligible cost for transitioning from LoRA to PiSSA.
@@
Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA. In this visualization, blue modules represent parts of the model whose parameters are frozen during training, while orange modules indicate components that require updates. QLoRA quantizes the pretrained matrix in LoRA to 4-bit, whereas QPiSSA quantizes the residual matrix in PiSSA.
Table 1: Comparison of similarities and differences between PiSSA and LoRA. In this table, bold highlights the model’s primary component, while underline denotes the residual component.
LoRA PiSSA
Forward $Y=X(\textbf{W}+\underline{\Delta W})=X(\textbf{W}+\underline{AB})$ $Y=X(\underline{W^{res}}+\textbf{W}^{\textbf{pri}})=X(\underline{W^{\text{res}}}+\textbf{AB})$
$\underline{A}\sim\mathcal{N}(0,\sigma^{2})\in\mathbb{R}^{m\times r}$ $\textbf{A}=U_{[:,\textbf{:r}]}\,S_{[\textbf{:r},\textbf{:r}]}^{1/2}\in\mathbb{R}^{m\times r}$
Initialization $\underline{B}=0\in\mathbb{R}^{r\times n}$ $\textbf{B}=S_{[\textbf{:r},\textbf{:r}]}^{1/2}\,V_{[:,\textbf{:r}]}^{\top}\in\mathbb{R}^{r\times n}$
$\underline{W^{\text{res}}}=U_{[:,\underline{r:}]}\,S_{[\underline{r:},\underline{r:}]}\,V_{[:,\underline{r:}]}^{\top}\in\mathbb{R}^{m\times n}$
Gradient $\frac{\partial L}{\partial A}=X^{\top}\left(\frac{\partial L}{\partial Y}\right)\underline{B}^{\top}\rightarrow\underline{0}$ $\frac{\partial L}{\partial A}=X^{\top}\left(\frac{\partial L}{\partial Y}\right)\textbf{B}^{\top}\rightarrow\textbf{Principal}$
$\frac{\partial L}{\partial B}=\underline{A^{\top}}X^{\top}\left(\frac{\partial L}{\partial Y}\right)\rightarrow$ Random Direction $\frac{\partial L}{\partial B}=\textbf{A}^{\top}X^{\top}\left(\frac{\partial L}{\partial Y}\right)\rightarrow\textbf{Principal}$
Comparison Fine-tunes noise while freezing W. Fine-tunes principal parts freezing $W^{\text{res}}$.
Slow convergence and underperformance. Fast convergence and better performance.
QLoRA cannot reduce quantization error. QPiSSA can reduce quantization error.
1 Introduction
Fine-tuning large language models (LLMs) is a highly effective technique for boosting their capabilities in various tasks luo2023wizardmath ; yu2023metamath ; luo2023wizardcoder ; li2023starcoder , ensuring models to follow instructions ouyang2022training ; zheng2024judging ; xu2023wizardlm , and instilling models with desirable behaviors while eliminating undesirable ones bai2022training ; rafailov2024direct . However, the fine-tuning process for very large models is accompanied by prohibitive costs. For example, regular 16-bit fine-tuning of a LLaMA 65B parameter model requires over 780 GB of GPU memory dettmers2024qlora , and the VRAM consumption for training GPT-3 175B reaches 1.2TB hu2021lora . Consequently, various parameter-efficient fine-tuning (PEFT) xu2023parameter ; han2024parameter methods have been proposed to reduce the number of parameters and memory usage required for fine-tuning. Due to the ability to maintain the performance of full fine-tuning without adding additional inference latency, Low-Rank Adaptation (LoRA) hu2021lora has emerged as a popular PEFT method.
LoRA hu2021lora hypothesizes that the modifications to parameter matrices during fine-tuning exhibit low-rank properties. As depicted in Figure 1(b), for a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA substitutes the updates with a low-rank decomposition $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, and the rank $r\ll\text{min}(m,n)$. For $Y=XW$, the modified forward pass is as follows:
$Y=X(W+\Delta W)=X(W+AB),$ (1)
A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially. LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix $W$, instead optimizing the injected, significantly smaller low-rank matrices $A,B$. Thus, it could reduce the number of trainable parameters by 10,000$\times$ and the GPU memory requirement by 3$\times$ hu2021lora . LoRA is capable of achieving comparable performance to full parameter fine-tuning. By integrating the quantization of pre-trained matrices $W$, LoRA also enables reducing the average memory requirements by 16$\times$ dettmers2024qlora . Meanwhile, the adapters can still utilize higher precision weights, thus, the quantization usually does not significantly degrade the performance of LoRA.
According to Equation 1, the gradients of A and B are $\frac{\partial L}{\partial A}=X^{\top}\left(\frac{\partial L}{\partial Y}\right)B^{\top}$ and $\frac{\partial L}{\partial B}=A^{\top}X^{\top}\left(\frac{\partial L}{\partial Y}\right)$. Compared to full fine-tuning, using LoRA initially does not change the output $Y$ for the same input $X$, so the magnitude and direction of gradient are primarily determined by the values of $A$ and $B$. Since $A$ and $B$ are initialized with Gaussian noise and zeros in LoRA, the gradients could be small and uninformative for a long time, leading to slow convergence in the fine-tuning process. We also observe this phenomenon empirically, as LoRA often wastes much time around the initial point.
@@ 1 Introduction
Figure 2: We illustrate the two key advantages of PiSSA: converging faster and better, and reducing quantization error. In the left figure, we use a toy example to show PiSSA’s faster convergence, where we first train a two-layer MLP classifying odd numbers of MNIST, and then fine-tune the model on even numbers. PiSSA finds the right direction more quickly and achieves a lower loss with the same number of steps. In the right figure, PiSSA reduces quantization error more effectively than LoftQ li2023loftq , with an optional 5-iteration SVD for further error reduction, as detailed in Appendix E.
@@
Because the principal components $W^{pri}$ are preserved in the adapter at full precision, an additional benefit of PiSSA is that when applying quantization to the frozen part $W^{res}$, we can significantly reduce the quantization error compared to QLoRA (which quantizes the entire $W$), as illustrated in Figure 2(b). Therefore, PiSSA is even more compatible with quantization than LoRA, making it a superior plug-and-play substitution for LoRA.
Our paper makes several significant contributions:
  • We analyze the initial gradient magnitude and direction in LoRA, demonstrating that $A$ initially has a zero gradient and $B$ has a random gradient, which slows down convergence and may lead to convergence at suboptimal local minima.
  • We propose PiSSA initialization, a novel method that approximates the optimization direction of full-parameter fine-tuning by adapting a model’s principal components. To our knowledge, PiSSA is the first to apply SVD to the original model, using principal singular values and vectors to initialize the adapter for fine-tuning, while keeping the residual components frozen. Experiments show that PiSSA converges faster and outperforms LoRA.
  • We combine PiSSA with NF4 quantization to propose QPiSSA, which reduces quantization error by about 20% compared to QLoRA, while maintaining the fast convergence and high performance of PiSSA.
@@ 2 Related Works
The vast complexity and computational needs of large language models (LLMs) with billions of parameters present significant hurdles in adapting them for specific downstream tasks. Parameter Efficient Fine-Tuning (PEFT) xu2023parameter ; han2024parameter emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning. PEFT encompasses strategies like partial fine-tuning zaken2021bitfit ; lawton2023neural ; zhao2020masking ; sung2021training ; ansell2021composable ; xu2021raise ; guo2020parameter ; fu2023effectiveness , soft prompt fine-tuning  hambardzumyan2021warp ; lester2021power ; li2021prefix ; liu2023gpt ; vu2021spot ; asai2022attempt ; wang2023multitask , non-linear adapter fine-tuning houlsby2019parameter ; lin2020exploring ; lei2024conditional ; he2021towards ; ruckle2020adapterdrop ; zhao2022tiny ; pfeiffer2020adapterfusion ; he2023mera ; mahabadi2021parameter ; chronopoulou2023adaptersoup , and low rank adapter based fine-tuning li2018measuring ; aghajanyan2020intrinsic ; hu2021lora ; zhang2022adaptive .
LoRA hu2021lora injects trainable adapters to the linear layers. After fine-tuning, these adaptations can be re-parameterized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning. Following LoRA, AdaLoRA zhang2022adaptive dynamically learns the rank size needed for LoRA in each layer of the model. DeltaLoRA zi2023delta updates the original weights of the model using parameters from adapter layers, enhancing LoRA’s representational capacity. LoSparse li2023losparse incorporates LoRA to prevent pruning from eliminating too many expressive neurons. DoRA liu2024dora introduces a magnitude component to learn the scale of $\Delta W$ while utilizing the original AB as a direction component of $\Delta W$. Unlike LoRA and its successors, which focus on learning low-rank approximations of weight updates, our PiSSA directly tunes the essential low-rank parts of the model while keeping the noisier, high-rank, and nonessential parts frozen. Although our approach differs in philosophy from LoRA, it shares most of LoRA’s structural benefits and can be extended by these methods to enhance its performance.
QLoRA dettmers2024qlora integrates LoRA with 4-bit NormalFloat (NF4) quantization, along with Double Quantization and Paged Optimizers, enabling the fine-tuning of a 65B parameter model on a single 48GB GPU while preserving the performance of full 16-bit fine-tuning tasks. QA-LoRA xu2023qa introduces group-wise operators to increase the degree of freedom in low-bit quantization. LoftQ li2023loftq reduces quantization error by decomposing the quantization error matrix of QLoRA and retaining the principal components with an adapter. PiSSA can also be combined with quantization techniques, and we have found that PiSSA significantly reduces quantization error compared to QLoRA and LoftQ.
@@ 3 PiSSA: Principal Singular Values and Singular Vectors Adaptation
Similar to LoRA, the gradients of $A$ and $B$ are also given by $\frac{\partial L}{\partial A}=X^{\top}\left(\frac{\partial L}{\partial Y}\right)B^{\top}$ and $\frac{\partial L}{\partial B}=A^{\top}X^{\top}\left(\frac{\partial L}{\partial Y}\right)$. Since elements of $\mathbf{s}_{[:r]}$ $\gg$ elements of $\mathbf{s}_{[r:]}$, the trainable adapter $W^{pri}=AB$ contains the most essential directions of $W$. In the ideal case, training $AB$ mirrors the process of fine-tuning the entire model despite using fewer parameters. The ability to directly fine-tune the most essential part of a model enables PiSSA to converge faster and better. In contrast, LoRA initializes the adapters $A$ and $B$ with Gaussian noise and zeros while keeping $W$ frozen. Consequently, the gradients are small or in random directions during the early stages of fine-tuning, possibly introducing much waste of gradient descent steps. Moreover, an inferior initialization might lead to suboptimal local minimum, resulting in worse generalization performance.
Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits. These include but are not limited to the capability of fine-tuning a model with a reduced number of trainable parameters, quantizing the residual model to decrease memory consumption during forward propagation in training, and easy deployment. The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model. Employing the Fast SVD technique halko2011finding allowed PiSSA to finish initialization in several seconds (Appendix B), which is a negligible cost.
@@
4 QPiSSA: An Extension Method with Lower Quantization Error
Quantization divides the value range of a matrix into several continuous regions, and maps all values falling inside a region into the same “quantized” value. It is an effective technique to reduce the memory consumption of forward propagation. At the same time, LoRA greatly reduces the backward memory requirement, making it highly suitable to use LoRA and quantization together, where the base model is quantized for memory-efficient forward propagation, and the LoRA adaptors are kept in full precision for accurate backward parameter updates. One representative previous work, QLoRA, quantizes the base model to Normal Float 4-bit (NF4) and initializes the full-precision $A$ and $B$ with Gaussian-Zero initialization. Therefore, the overall error is given by:
@@ 4 QPiSSA: An Extension Method with Lower Quantization Error
where $||M||_{*}$ denotes the nuclear norm (also known as the trace norm fan1951maximum ), defined as:
@@
Since the residual model has removed the large-singular-value components, $W^{res}$ has a narrower distribution than that of $W$, as can be seen in Figures 3(a) and 3(b) (comparing the singular value distributions of $W$ and $W^{res}$), as well as Figures 3(c) and 3(f) (comparing the value distributions of $W$ and $W^{res}$), which is beneficial for reducing the quantization error. Moreover, given that NF4 is optimized for data with a normal distribution, as discussed by Dettmers et al. dettmers2024qlora , we fit the values of $W$ and $W^{res}$ to a Gaussian distribution respectively. As illustrated in Figures 3(c) and 3(f), $W^{res}$ aligns more closely with a Gaussian distribution and exhibits a smaller standard deviation, making it more suitable for applying NF4 than $W$. Both the above lead QPiSSA to achieve a significantly lower quantization error than QLoRA, shown in Figures 3(d) and 3(e).
@@
Figure 3: Visualizations of LLaMA 2-7B’s “layers[0].self_attn.q_proj” matrix, with distributions for the full model shown in Appendix G. Figures (a), (b), (d), and (e) display the singular values of $W$, $W^{res}$, $W-nf4(W)$, and $W^{res}-nf4(W^{res})$, respectively. Figures (c) and (f) show the data distributions of $W$ and $W^{res}$.